# Always skip attention

### v2

@@ 5.1 Singular Value Decomposition(SVD)
Singular Value Decomposition of a matrix $\mathbf{X}\,{\in}\,\mathbb{R}^{n{\times}d}$ is the factorization of $\mathbf{X}$ into the product of three matrices $\mathbf{X=U\Sigma V^{T}}$ where the columns of $\mathbf{U}\,{\in}\,\mathbb{R}^{n{\times}n}$ and $\mathbf{V}\,{\in}\,\mathbb{R}^{d{\times}d}$ are orthonormal and $\mathbf{\Sigma}\,{\in}\,\mathbb{R}^{n{\times}d}$ is a rectangular diagonal matrix whose diagonal elements are non-negative in descending order and represent the singular values such that $\Sigma_{ii}\,{=}\,\sigma_{i}$. As observed in Eq. 4, we can reduce the condition number by increasing the minimal singular value while keeping the maximal singular value unchanged. Hence, our goal is to reconstruct $\mathbf{\tilde{X}}=\mathbf{U\tilde{\Sigma}}\mathbf{V^{T}}$, where $\mathbf{\tilde{\Sigma}}$ is formed by amplifying non-maximal elements of $\mathbf{\Sigma}$ while keeping the maximal element unchanged. Our SVD token graying pipeline is presented in Algorithm 1.
However, directly applying SVD to matrices is computationally expensive, with a cost of $O(nd\cdot\text{min}(n,d))$ leading to longer training time. Tab. 1 demonstrates that training a vision transformer is significantly slower when using SVD reconstruction. Therefore, in the next subsection, we introduce a more efficient token graying method using the discrete cosine transform.
@@ 5.2 Discrete Cosine Transform(DCT)
There are several variants of the DCT and in this work we choose DCT-II which is probably the most commonly used form. Given a real-valued sequence data $\mathbf{x}\in\mathbb{R}^{N\times 1}$, the 1D DCT sequence $\hat{\mathbf{x}}$ is defined as:
$\hat{\mathbf{x}}_{k}=\alpha_{k}\sum_{i=0}^{N-1}$ $x_{i}\cos\left[\frac{\pi(2i+1)k}{2N}\right],\text{for }k=0,1,...,N-1$ (8)
where $\text{where }\alpha_{k}=\begin{cases}{\sqrt{\frac{1}{N}}},\text{if }k=0\\ {\sqrt{\frac{2}{N}}},\text{if }k\neq 0\\ \end{cases}$
Simplicitly, 1D DCT is a linear transformation, Eq. 8 can be written with the formula $DCT(\mathbf{x})\,{=}\,\mathbf{\hat{x}}\,{=}\,D\mathbf{x}$, where $D_{i,k}$ is the 1D DCT basis and orthogonal.
@@
For the reconstruction process, since the DCT expresses a signal as a sum of cosine functions with different frequencies, the IDCT combines these frequency components to recover the original signal. Hence, IDCT$(\hat{x})\,{=}\,\tilde{\mathbf{x}}=D^{T}\mathbf{\hat{x}}$. In our case, we have a token matrix $\mathbf{X}\,{\in}\,\mathbb{R}^{n{\times}d}$ and we generalize the 1D DCT to a 2D DCT using the formula $\mathbf{\hat{X}}\,{=}\,D\mathbf{X}D^{T}$. In natural images, the dominant singular vector—corresponding to the largest singular value—typically captures the low frequency content of the image. Intuitively, this observation suggests that DCT reconstruction can serve as an approximation for SVD reconstruction. Our DCT token graying pipeline is presented in Algorithm 2.
DCT avoids the expensive computational cost of SVD and has a complexity of $O(nd\cdot\log(nd))$. Tab. 1 shows that training ViT using the DCT algorithm is significantly faster than SVD algorithm.
@@
1$\triangleright$ Input: training data, amplification coefficient $\epsilon\in(0,1]$
2$\triangleright$ Training:
@@
5:     $D\mathbf{\hat{X}}D^{T}\leftarrow\mathbf{X}$triangleright Compute DCT
6:     $\mathbf{\hat{X}}_{\text{norm}}\leftarrow|\mathbf{\hat{X}}|/\text{max}(|\mathbf{\hat{X}}|)$triangleright Normalize elements $\in(0,1]$
7:     $\mathbf{\hat{X}}\leftarrow\mathbf{\hat{X}}_{\text{norm}}^{\epsilon}\cdot\text{sign}(\mathbf{\hat{X}})\cdot\text{max}(|\mathbf{\hat{X}}|)$triangleright Amplify
8:     $\mathbf{\tilde{X}}\leftarrow\hat{D}^{T}\mathbf{\hat{X}}\hat{D}$triangleright Inverse DCT
9:     Patch Embedding: $\mathbf{Z}\leftarrow\text{PatchEmbed}(\tilde{\mathbf{X}})$
@@ 6 Experiments
[…]
@@ 6.1 Supervised learning ViT
Method Top-1 Acc Top-5 Acc
$\frac{1}{k}(\mathbf{QK})$ 43.0 % 62.7 %
$\frac{1}{k}(\mathbf{QK})$ + SVD 44.3 % 68.8 %
$\frac{1}{k}(\mathbf{QK})$ + DCT 44.2 % 68.8 %
$\operatorname*{softmax}(\mathbf{QK})$ 43.0 % 67.3 %
$\operatorname*{softmax}(\mathbf{QK})$ + SVD 44.7 % 69.7 %
$\operatorname*{softmax}(\mathbf{QK})$ + DCT 44.8 % 69.7 %
@@
Results. In Tab. 2, we first demonstrate the results for the linear case and the regular $\operatorname*{softmax}$ case using our methods on a small model (ViT-Tiny) and the small-scale Tiny-ImageNet dataset. Given that Theorem Proposition 4.2 and Proposition Eq. 5 are mathematically explained using linear self-attention, we train the ViT-Tiny model with linear self-attention, which aligns well with our analysis in Fig. 4. Furthermore, as shown in Fig. 5, both SVDTG and DCTTG lead to better conditioning of SAB output embeddings throughout training epochs. Next, we train the model on a larger dataset using the default $\operatorname*{softmax}$ self-attention, further validating that our approach generalizes to $\operatorname*{softmax}$ self-attention. Moreover, SVDTG is the most direct method to improve the conditioning of a matrix, and as demonstrated in Tab. 3, it leads to notable performance gains. To address the high computational cost—approximately six times slower than regular ViT training—we approximate SVDTG using DCTTG, which achieves better token embeddings conditioning and yields comparable results too. As shown in Fig. 6, our DCTTG method produces condition numbers across all layers that indicate better token conditioning in almost every layer. Finally, we evaluate our methods on different variants of ViTs trained on ImageNet-1K datasets, where our DCTTG method demonstrates superior performance compared to vanilla models in terms of Top-1 and Top-5 classification accuracy, while also achieving better conditioning of SAB output embeddings across all layers in the ViT-B model.
@@
Table 4: Top-1 and Top-5 classification accuracy on ImageNet-1K dataset using DCT token graying on different ViTs
@@ 6.2 Self-supervised learning ViT
In this subsection, we validate our methods on self-supervised learning models as well as on fine-tuning pretrained models. Self-supervised learning is a powerful technique in which models learn representations directly from unlabeled data. Masked Image Modeling (MIM) [9] learns representations by reconstructing images that have been corrupted through masking.
@@
Finetuning and reuslts. In the finetuning stage, we fine-tune the pre-trained models on ImageNet-1k for classification tasks using cross-entropy loss. We use a batch size of $512$, a base learning rate of $5\times 10^{-4}$, a weight decay of $0.05$, and train for 100 epochs. As shown in Tab. 5, our DCTTG method outperforms the vanilla MAE model by 0.2% in both top-1 and top-5 classification accuracy.
@@ 8 Conclusion
This paper provides a theoretical demonstration that the output embeddings of linear self-attention blocks—when lacking skip connections—are inherently ill-conditioned, a finding that empirically extends to conventional $\operatorname*{softmax}$ self-attention blocks as well. Furthermore, we show that skip connections serve as a powerful regularizer by significantly improving the conditioning of self-attention blocks’ output embeddings. In addition, we use the condition of the self-attention output embeddings as a proxy for that of its Jacobian, empirically demonstrating that input tokens with improved conditioning lead to a better-conditioned Jacobian. Building on this insight, we introduce SVDTG and DCTTG, simple yet effective methods designed to improve the condition of input tokens. We then validate their effectiveness on both supervised and unsupervised learning tasks. Ultimately, we hope these insights will guide the vision community in designing more effective and efficient ViT architectures for downstream applications.
Supplementary Material
@@ Appendix A Theoretical framework
In this section we give the proof of our Eq. 5.
@@
$\kappa(\mathbf{XW_{\text{Q}}})\leq\kappa(\mathbf{X})\cdot\kappa(\mathbf{W_{\text{Q}}}).\$ (11)
@@
$\kappa(\mathbf{XW_{\text{Q}}})\leq C_{\text{Q}}\cdot\kappa(\mathbf{X}).\$ (12)
@@
Next, we give the proof of our Proposition 4.2
@@
This completes our proof for Proposition 4.2.
@@ Appendix B ConvMixer
In Fig. 2, we demonstrate that the performance of the ConvMixer Tiny model does not degrade when skip connections are removed. This observation contrasts with Vision Transformers (ViTs), where the absence of skip connections in the self-attention mechanism leads to a noticeable performance drop.
@@ Appendix C Empirical Neural Tangent Kernel
The Neural Tangent Kernel (NTK) describes the evolution of deep neural networks during training by gradient descent. In this paper, we propose to measure the condition of the self-attention output embeddings as a proxy for its Jacobian condition, since analyzing the transformer model NTK is computationally expensive and unrealistic. However, in Fig. 7 and Fig. 8, we empirically demonstrate the Jacobian of the self-attention mechanism using SVDTG and DCTTG. Intrinsically, the self-attention exhibits a disproportionately ill-conditioned spectrum, and using our TG methods, we observe an improvement in this regard.

### v3

@@ 5.1 Singular Value Decomposition (SVD)
Singular Value Decomposition of a matrix $\mathbf{X}\,{\in}\,\mathbb{R}^{n{\times}d}$ is the factorization of $\mathbf{X}$ into the product of three matrices $\mathbf{X=U\Sigma V^{T}}$ where the columns of $\mathbf{U}\,{\in}\,\mathbb{R}^{n{\times}n}$ and $\mathbf{V}\,{\in}\,\mathbb{R}^{d{\times}d}$ are orthonormal and $\mathbf{\Sigma}\,{\in}\,\mathbb{R}^{n{\times}d}$ is a rectangular diagonal matrix whose diagonal elements are non-negative in descending order and represent the singular values such that $\Sigma_{ii}\,{=}\,\sigma_{i}$. As observed in Eq.˜4, we can reduce the condition number by increasing the minimal singular value while keeping the maximal singular value unchanged. Hence, our goal is to reconstruct $\mathbf{\tilde{X}}=\mathbf{U\tilde{\Sigma}}\mathbf{V^{T}}$, where $\mathbf{\tilde{\Sigma}}$ is formed by amplifying non-maximal elements of $\mathbf{\Sigma}$ while keeping the maximal element unchanged. Our SVD token graying pipeline is presented in Algorithm˜1.
However, directly applying SVD to matrices is computationally expensive, with a cost of $O(nd\cdot\text{min}(n,d))$ leading to longer training time. Tab.˜1 demonstrates that training a ViT is significantly slower when using SVD reconstruction. Therefore, in the next subsection, we introduce a more efficient Token Graying (TG) method using the discrete cosine transform.
@@ 5.2 Discrete Cosine Transform (DCT)
There are several variants of the DCT and in this work we choose DCT-II which is one of the most commonly used forms. Given a real-valued sequence data $\mathbf{x}\in\mathbb{R}^{N\times 1}$, the 1D DCT sequence $\hat{\mathbf{x}}$ is defined as:
$\hat{\mathbf{x}}_{k}=\alpha_{k}\sum_{i=0}^{N-1}$ $x_{i}\cos\left[\frac{\pi(2i+1)k}{2N}\right],\text{for }k=0,1,.,N-1$ (8)
[…]
Simplicitly, 1D DCT is a linear transformation, Eq.˜8 can be written with the formula $DCT(\mathbf{x})\,{=}\,\mathbf{\hat{x}}\,{=}\,D\mathbf{x}$, where $D_{i,k}$ is the 1D DCT basis and orthogonal.
@@
[…]
DCT avoids the expensive computational cost of SVD and has a complexity of $O(nd\cdot\log(nd))$. Tab.˜1 shows that training ViT using the DCT algorithm is significantly faster than SVD algorithm.
@@
[…]
@@
5:  $D\mathbf{\hat{X}}D^{T}\leftarrow\mathbf{X}$ $\triangleright$ Compute DCT
[…]
8:  $\mathbf{\tilde{X}}\leftarrow\hat{D}^{T}\mathbf{\hat{X}}\hat{D} \triangleright$ Inverse DCT
[…]
@@ 6 Experiments
Vision Transformers (ViTs) have emerged as powerful models in the field of computer vision, demonstrating remarkable performance across a variety of tasks. This section is dedicated to validating and analyzing our proposed algorithm for both supervised and self-supervised learning in ViTs. For all experiments, unless specified otherwise, we use $\epsilon\,{=}\,0.95$.
@@ 6.1 Supervised learning ViT
Method Top-1 Acc (%) Top-5 Acc (%)
$\frac{1}{k}(\mathbf{QK})$ 43.0 62.7
$\frac{1}{k}(\mathbf{QK})$ + SVD 44.3 68.8
$\frac{1}{k}(\mathbf{QK})$ + DCT 44.2 68.8
$\operatorname*{softmax}(\mathbf{QK})$ 43.0 67.3
$\operatorname*{softmax}(\mathbf{QK})$ + SVD 44.7 69.7
$\operatorname*{softmax}(\mathbf{QK})$ + DCT 44.8 69.7
@@
Results. In Tab.˜2, we first demonstrate the results for the linear case and the regular $\operatorname*{softmax}$ case using our methods on a small model (ViT-Tiny) and the small-scale Tiny-ImageNet dataset. Given that Theorem Proposition˜4.2 and Proposition Eq.˜5 are mathematically explained using linear self-attention, we train the ViT-Tiny model with linear self-attention, which aligns well with our analysis in Fig.˜3. Furthermore, as shown in Fig.˜4, both SVDTG and DCTTG lead to better conditioning of SAB output embeddings throughout training epochs. Next, we train the model on a larger dataset using the default $\operatorname*{softmax}$ self-attention, further validating that our approach generalizes to $\operatorname*{softmax}$ self-attention. Moreover, SVDTG is the most direct method to improve the conditioning of a matrix, and as demonstrated in Tab.˜3, it leads to notable performance gains. To address the high computational cost—approximately six times slower than regular ViT training—we approximate SVDTG using DCTTG, which achieves better token embeddings conditioning and yields comparable results too. As shown in Fig.˜5, our DCTTG method produces condition numbers across all layers that indicate better token conditioning in almost every layer. Finally, we evaluate our methods on different variants of ViTs trained on ImageNet-1K datasets, where our DCTTG method demonstrates superior performance compared to vanilla models in terms of Top-1 and Top-5 classification accuracy, while also achieving better conditioning of SAB output embeddings across all layers in the ViT-B model.
@@
Table 4: Top-1 and Top-5 classification accuracy on ImageNet-1K dataset using DCT token graying on different ViTs configurations.
@@ 6.2 Self-supervised learning ViT
In this subsection, we validate our methods on self-supervised learning models as well as on fine-tuning pretrained models [12, 7]. Self-supervised learning is a powerful technique in which models learn representations directly from unlabeled data. Masked Image Modeling (MIM) [12] learns representations by reconstructing images that have been corrupted through masking.
@@
Finetuning and reuslts. In the finetuning stage, we fine-tune the pre-trained models on ImageNet-1k for classification tasks using cross-entropy loss. We use a batch size of $512$, a base learning rate of $5\times 10^{-4}$, a weight decay of $0.05$, and train for 100 epochs. As shown in Tab.˜5, our DCTTG method outperforms the vanilla MAE model by 0.2% in both top-1 and top-5 classification accuracy.
@@ 8 Conclusion
This paper provides a theoretical demonstration that the output embeddings of linear self-attention blocks, when lacking skip connections, are inherently ill-conditioned, a finding that empirically extends to conventional $\operatorname*{softmax}$ self-attention blocks as well. Furthermore, we show that skip connections serve as a powerful regularizer by significantly improving the conditioning of self-attention blocks’ output embeddings. In addition, we use the condition of the self-attention output embeddings as a proxy for that of its Jacobian, empirically demonstrating that input tokens with improved conditioning lead to a better-conditioned Jacobian. Building on this insight, we introduce SVDTG and DCTTG, simple yet effective methods designed to improve the condition of input tokens. We then validate their effectiveness on both supervised and unsupervised learning tasks. Ultimately, we hope these insights will guide the vision community in designing more effective and efficient ViT architectures for downstream applications.
Acknowledgements. This work was supported by the Australian Institute for Machine Learning (AIML) and the CSIRO’s Data61 Embodied AI Cluster.
@@ Appendix A Theoretical framework
In this section, we give the proof of our Eq.˜5.
@@
$\kappa(\mathbf{XW_{\text{Q}}})\leq\kappa(\mathbf{X})\cdot\kappa(\mathbf{W_{\text{Q}}}).\\$ (11)
@@
$\kappa(\mathbf{XW_{\text{Q}}})\leq C_{\text{Q}}\cdot\kappa(\mathbf{X}).\\$ (12)
@@
Next, we give the proof of our Proposition˜4.2
@@
This completes our proof for Proposition˜4.2.
@@ Appendix B ConvMixer
In Fig.˜6, we demonstrate that the performance of the ConvMixer Tiny model does not degrade when skip connections are removed. This observation contrasts with Vision Transformers (ViTs), where the absence of skip connections in the self-attention mechanism leads to a noticeable performance drop.
@@ Appendix C Empirical Neural Tangent Kernel
The Neural Tangent Kernel (NTK) describes the evolution of deep neural networks during training by gradient descent. In this paper, we propose to measure the condition of the self-attention output embeddings as a proxy for its Jacobian condition, since analyzing the transformer model NTK is computationally expensive and unrealistic. However, in Fig.˜7 and Fig.˜8, we empirically demonstrate the Jacobian of the self-attention mechanism using SVDTG and DCTTG. Intrinsically, the self-attention exhibits a disproportionately ill-conditioned spectrum, and using our TG methods, we observe an improvement in this regard.